# Advaith Sethuraman, 2505 Hayward St., Ann Arbor, MI 48109

### v1

@@ 3.4 AUV Survey Mission Planning
The Iver3 AUV has two standard patterns for pre-planned missions: lawnmower (LM) and object identification (OID). The LM pattern is a common down-and-back structure, as depicted by the dotted lines in Fig. 5b. In one leg of a LM survey, there are three stages: descent, survey, and ascent. The AUV must ascend, turn around, and then descend before collecting sonar data for each leg in order to geolocalize itself, as global positioning system (GPS) data is not available underwater. The leg length ($l$) in Fig. 5b denotes the distance that the AUV was underwater collecting SSS data, and the half-swath width ($\frac{\lambda}{2}$) describes the distance on the seafloor covered by the fan-shaped sonar beams for one of the transducer arrays. $w_{s}$ is the total survey width of a planned LM mission, and $\delta_{w}$ is the width of each individual leg. The AUV also takes in either a depth from surface ($d_{s}$) or a height from bottom ($h_{b}$) parameter which describes the altitude it will descend to before collecting SSS data.
@@ 3.5 Post-Processing
The Iver3 AUV stores each leg of a sonar survey separately as a .JSF file (proprietary format). These files are readable by the free Edgetech Discover software. The software allows for limited sonar image processing and exporting to .PNG. The Discover software normalizes the sonar image intensity and applies Time Varying Gain (TVG). As sound moves through water, absorption loss and spreading cause decreased signal strength. As a result, the returned echos from objects will also have reduced intensity. TVG addresses this shortcoming and ensures the intensity of the return on the sonar image is not range-dependent MacLennan (1986).
Pixel-wise labeling of the post-processed sonar images was conducted by a team of three researchers with shared labeling guidelines. The labels were then reviewed in-depth by a marine archaeologist from the State of Michigan who is an expert on the shipwrecks at TBNMS. There are two labels: “shipwreck” and “other.” “Shipwreck” consists of the primary wrecks as well as any debris. If part of the ship is obscured in an acoustic shadow, the label was extrapolated into the shadowed region to follow the expected shape of the wreck based off of expert knowledge. The labels are exported to a standard binary mask format where 0 represents the “other” class and 1 represents the “shipwreck” class.
As survey lengths ($l$) can vary for each mission, the resulting sonar images are of different lengths. The dataset provides full-sized images, but we square-cropped these images to be input to deep neural networks. Crops are generated by padding the full-size images from $h\times w$ to $H\times w$ according to
$H=h+s-[(h-w)\%s]$ (1)
where $s=100$ is the stride length in pixels. Then, images of dimensions $w\times w$ are cropped every $s$ pixels down the length of the padded image ($H$). A visualization of this process is shown in Fig. 7. The only other data processing step applied to the images is a normalization around 0 based on the mean and standard deviation of the pixel values of the entire dataset.
Figure 7: Visualization of the strided cropping of the original side scan sonar image to process before inputting into neural networks. Each colored square represents a $w\times w$ crop of the image, with a stride length of $s=100\ px$.
@@ 4 Dataset Organization
The dataset is organized by the SSS image and their corresponding labels. We provide the labels as segmentation masks of objects belonging to a shipwreck in the SSS image, including the shipwreck itself and any debris associated with the shipwreck. This includes parts of the shipwreck that may have come off from the shipwreck and parts of the ship that has been dislodged. The images and labels are separated into test and train sets using a 50/50 split. We select a 50/50 split to ensure that the test set is large enough to capture the variance of the shipwrecks sites, ensuring that the robustness of deep learning models can be evaluated through our dataset. In addition, the images are grouped on a per-site basis. This is done to ensure that SSS images that see the same terrain, shipwreck, or debris are not shared between the test and train sets of our dataset.
@@ 4.1 Released Dataset File Structure
In Fig. 8, the released data directory structure is shown. The data is split into test and train, according to our iterative approach as described in this paper. Within the test and train directories, the sonar images and their corresponding binary labels are stored. Each sonar image (.PNG format) corresponds to one leg of a survey mission. The site of the survey is indicated in the image file names.
Table 1: Details of site characteristics, image quality, and number of images per site. Condition is in increasing order with 8 corresponding to most intact. SSS quality is in increasing order with 3 corresponding to highest quality. Terrain corresponds to categorical terrain types. ${}^{*{\dagger},{\ddagger}}$ denote shipwrecks that were imaged in a single survey and thus considered as a single “site.”
Name
Depth (m)
Length (m) Beam (m) Material Condition SSS Quality Terrain # Imgs.
Training
Alpena ${}^{{\dagger}}$
$4$ $2$ C 10
Bay ${}^{{\dagger}}$
$3.4$
$44.5$ $8.8$ Wood $4$ $2$ C 10
D.M. Wilson
$12.2$
$54.6$ $9.8$ Wood $6$ $3$ B 22
D.R. Hanna
$41.1$
$162.2$ $17.1$ Steel $8$ $2$ A 5
E.B. Allen
$30.5$
$40.8$ $7.9$ Wood $6$ $3$ D 24
Egyptian
$79.2$
$70.7$ $11.0$ Wood $5$ $2$ D 15
Grecian
$30.5$
$90.2$ $12.2$ Steel $7$ $2$ B 5
Harvey ${}^{{\dagger}}$
$4.6$
$49.4$ $10.1$ Wood $4$ $2$ C 10
Heart Failure
$5.5$
Wood $2$ $1$ B 12
Isaac M. Scott
$53.3$
$159.7$ $16.5$ Steel $8$ $2$ C 6
Montana
$19.2$
$71.9$ $11.0$ Wood $6$ $3$ A 8
Oscar T. Flint
$9.1$
$66.4$ $11.3$ Wood $5$ $2$ A 22
Pewabic
$50.3$
$61.0$ $9.4$ Wood $6$ $2$ A 5
W.P. Rend
$5.2$
$87.5$ $12.2$ Wood $5$ $3$ A 11
Testing
Artificial Reef
$3.0$
$1$ $1$ B 6
Barge No. 1
$12.8$
$94.2$ $13.4$ Wood $5$ $2$ A 15
B. ${}^{{\ddagger}}$
$4.6$
$41.1$ $5.8$ Wood $2$ $1$ B 10
Corsair
$55.5$
$40.8$ $7.3$ Wood $5$ $2$ C 4
Corsican
$48.8$
$34.1$ $7.6$ Wood $5$ $1$ C 6
James ${}^{{\ddagger}}$
$10.7$
$30.5$ $9.1$ Wood $4$ $1$ A 10
John F. $*{}^{*}$
$2.7$
$38.4$ $7.9$ Wood $3$ $1$ B 6
W.H. Gilbert
$77.7$
$100.0$ $12.8$ Steel $8$ $3$ C 6
Haltiner Barge
$5.2$
$24.4$ $10.1$ Wood $3$ $2$ A 13
L. Van Valkenburg
$18.3$
$39.0$ $7.9$ Wood $5$ $3$ B 19
${}^{{\ddagger}}$
$5.5$
$48.8$ $9.1$ Wood $5$ $1$ B 10
Monrovia
$42.7$
$136.6$ $17.1$ Steel $8$ $3$ A 8
$3.4$
$44.5$ $9.1$ Wood $2$ $1$ B 6
W.P. Thew
$25.6$
$40.2$ $7.3$ Wood $5$ $3$ A 17
Viator
$57.3$
$70.7$ $10.1$ Steel $8$ $2$ C 11
Table 1: Details of site characteristics, image quality, and number of images per site. Condition is in increasing order with 8 corresponding to most intact. SSS quality is in increasing order with 3 corresponding to highest quality. Terrain corresponds to categorical terrain types. ${}^{*{\dagger},{\ddagger}}$ denote shipwrecks that were imaged in a single survey and thus considered as a single “site.”
4.2 Test-Train Split Selection
Our goal with the test-train split is to ensure that each set provides enough diversity of data for both the training and evaluation of deep learning models. We inform our split selection through both expert-informed and data-driven metrics, with the target of capturing the underlying distribution of the data as uniformly as possible. To inform the test-train split selection, we group the sites based on three major categories: shipwreck condition, sonar image quality (SIQ), and terrain type. For each site, a vector score of length three is assigned based on these quantitative and qualitative metrics. A 50/50 test-train split is then performed using an iterative stratification on our multi-labeled data Sechidis et al. (2011).
@@ 4.2.1 Wreck Condition:
We leverage domain expert knowledge from scientists at TBNMS in order to assign wreck condition labels. Experts who have conducted dive surveys to inspect these ships provided a scale of 1-8 for wreck condition: 8 = complete and intact, 7 = complete but collapsed, 6 = semi-complete, 5 = partially intact, 4 = fragmented, 3 = fragmented and disarticulated, 2 = widely dispersed debris fields, and 1 = non-shipwreck cultural materials (e.g., artificial reef).
@@ 4.2.2 Sonar Image Quality Assessment:
The sonar image quality (SIQ) assessment is done based on well-developed metrics for optical image quality assessment (IQA). We utilize the no-reference, completely blind IQA metric Natural Image Quality Evaluator (NIQE) Mittal et al. (2013), which constructs a quality score from a natural scene statistic model trained on undistorted images. There are many challenges in evaluating the quality of a sonar image, the first of which being that many IQA metrics are targeted at measuring quality deterioration from compression. However, our sonar images are not compressed—rather, any distortion in the images is a result of when the Iver3 is surveying shallower sites, making it more vulnerable to drift from waves (Fig. 6d). If the shipwreck partially falls into the nadir, this can cause significant interference and acoustic shadows that disrupt the geometry of the wreck (Fig. 6e). Additionally, if the shipwreck falls into the outer edges of the sonar image, the increased noise of more distant acoustic returns causes significant noise in the final sonar image (Fig. 6f). Informed by the NIQE score, we assign classes of 1-3 to the images where 1 indicates poorer sonar image quality and 3 indicates higher image quality.

### v2

@@ 3.5 AUV Survey Mission Planning
The Iver3 AUV has two standard patterns for pre-planned missions: lawnmower (LM) and object identification (OID). The LM pattern is a common down-and-back structure, as depicted by the dotted lines in Fig. 5b. In one leg of a LM survey, there are three stages: descent, survey, and ascent. The AUV must ascend, turn around, and then descend before collecting sonar data for each leg in order to geolocalize itself, as global positioning system (GPS) data is not available underwater. The survey occurs at either a constant height from bottom ($h_{b}$) or depth from surface ($d_{s}$), and this parameter may change from site to site based on the depth of the wreck. Given that the maximum range of the DVL is 81m, and the maximum depth of the shipwreck sites surveyed was 79.2m, we may conclude that bottom-tracking was not lost during repeated descent/ascent stages. The leg length ($l$) in Fig. 5b denotes the distance that the AUV was underwater collecting SSS data, and the half-swath width ($\frac{s}{2}$) describes the distance on the seafloor covered by the sonar beams from one of the transducer arrays. For a planned LM mission, $w_{s}$ is the total survey width and $\delta_{w}$ is the width of each individual leg.
Table 3: Detailed survey parameter information including sonar range, leg length, survey depth (from surface) or survey height (from bottom). Note that sonar range and leg length are related to the width and height of a captured SSS image in meters respectively. All values were recorded prior to deployment of the AUV and cross-referenced with information from the .JSF files. Note we are describing each individual survey in this table rather than aggregated per site, as multiple surveys at one location may have had different parameters.
Survey
2*Sonar
Range (m)
Leg
Dist. (m)
Depth
Height
$R_{x}$
Range (cm)
$R_{y}$
20220523-141805–EBAllenLM 60 110 25 13.61 1.16
20220523-154917–EBAllenLM15mDFS 60 110 15 13.61 1.15
20220523-185512–Thew15DFS 60 150 15 13.61 1.17
20220523-204910–Thew15DFS54oa 60 150 15 13.61 1.17
20220524-134748–VanValkenburgLM10mDFS 60 100 10 13.61 1.17
20220524-150629–VanValkenburgOID10mDFS 60 150 10 13.61 1.13
20220524-161604–WilsonOID2mDFS 60 150 2 13.61 1.13
20220524-174036–WilsonBroadSearch2mDFS 60 1000 2 13.61 1.13
20220524-201003–MontanaOID8mDFS 60 150 8 13.61 1.13
20220527-134851–ViatorOID12mDFS 120 400 12 27.23 1.17
20220527-180547–MonroviaOID20mDFS 250 400 20 56.72 1.17
20220527-211838–HaltinerLM2mDFS 250 150 2 56.72 1.15
20220531-132042–Flint 120 400 2.5 27.23 1.15
20220531-145419–Flint_terrain 240 400 2.5 54.45 1.16
20220531-162857–Flint_120m 240 400 2.5 54.45 1.16
20220531-180338–Heart_Failure 260 200 1 58.99 1.17
20220601-143442–BargeNo1 250 600 2 56.72 1.16
20220601-171552–Rend 120 200 1 27.23 1.16
20220601-185052–Heart_Failure2 120 200 2 27.23 1.13
20220602-131155–Grecian 260 250 5 58.99 1.15
20220602-141250–Pewabic 240 400 22 54.45 1.15
20230605-141930–Egyptian1_100m_DFS45m_LM 200 400 45 45.38 1.14
20230605-173643–Egyptian2_130m_DFS32m_OID 260 300 32 58.99 1.14
20230605-193324–Gilbert1_130m_DFS15m_LM 260 200 15 58.99 1.13
20230605-202803–Gilbert2_160m_DFS25m_LM 320 400 25 72.61 1.13
20230606-181949–NellieGardner1_HFB_LM 120 400 3 27.23 1.13
20230606-195902–NellieGardner2_HFB_LM 70 400 3 15.88 1.13
20230607-122726–Monohansett1_HFB_LM 100 800 3 22.69 1.13
20230607-144904–MischelleyReef1_3mDFS_LM 80 500 3 18.15 1.13
20230607-164552–MischelleyReef2_3mDFS_LM 80 400 3 18.15 1.13
20230607-175515–NearShore1_2mHFB_LM 60 500 3 13.61 1.14
20230608-125411–Hanna1_100m_DFS20m_LM 300 600 10 68.07 1.14
20230608-154854–Terrain1_3mDFS_LM 160 2000 2.5 36.30 1.14
20230609-130155–Corsican1_100m_DFS5m_LM 300 400 5 68.07 1.14
20230609-155043–Corsican2_100m_DFS5m_LM 300 700 5 68.07 1.14
20230609-173705–Corsair1_700m_DFS12m_LM 300 500 12 68.07 1.14
20230609-194421–Exploratory1_15mDFS_LM 160 2000 15 36.30 1.14
20230614-121856–Rend 80 400 1 18.15 1.16
20230614-155731–Monohansett1_HFB_LM 70 300 3 15.88 1.14
20230614-173130–Davidson 70 420 1 15.88 1.14
20230615-124230–Scott1 300 600 15 68.07 1.16
20230615-140855–Scott2 300 400 10 68.07 1.16
20230615-152709–Wilson-2mDFS 120 600 3 27.23 1.16
20230615-165126–VanValkenburg3mDFS 240 600 3 54.45 1.17
20230616-125457–Shamrock1_1m_DFS 180 300 1 40.84 1.16
20230616-133407–Shamrock2_1m_DFS_2 140 300 1 31.76 1.16
20230616-143922–BargeNo1_23 220 600 2 49.92 1.16
20230616-155616–BargeNo1_Terrain 200 600 3 45.38 1.16
20230616-173938–Haltiner_Terrain 70 600 1.5 15.88 1.16
20230616-193812–Haltiner_Bilge 100 75 1.5 22.69 1.16
@@ 3.6 Post-Processing
The Iver3 AUV stores each leg of a sonar survey separately as a .JSF file, which is a proprietary format (EdgeTech JSF Spec., 2023). These files are readable by the free Edgetech Discover software. We used the Discover software for limited sonar image processing and exporting to .PNG. The Discover software normalizes the sonar image intensity and applies Time Varying Gain (TVG). As sound moves through water, absorption loss and spreading cause decreased signal strength. As a result, the returned echoes from objects will also have reduced intensity. TVG addresses this shortcoming and ensures the intensity of the return on the sonar image is not range-dependent (MacLennan, 1986). We do not perform slant-range correction within the Discover software (Chang et al., 2010). No offline post-processing of the .JSF data was performed using GPS information.
Pixel-wise labeling of the post-processed sonar images was conducted by a team of three researchers with shared labeling guidelines. The labels were then reviewed in-depth by a marine archaeologist from the State of Michigan who is an expert on the shipwrecks at TBNMS. There are two labels: “shipwreck" and “other." “Shipwreck" consists of the primary wrecks as well as any debris. If part of the ship is obscured in an acoustic shadow, the label was extrapolated into the shadowed region to follow the expected shape of the wreck based off of expert knowledge. The labels are exported to a standard binary mask format where 0 represents the “other" class and 1 represents the “shipwreck" class.
As survey lengths ($l$, in meters) can vary for each mission, the resulting sonar images are of different heights ($h$, in pixels). The dataset provides full-sized images of dimension $h\times w$, where $w$ is image width in pixels. We square-cropped these images into overlapping tiles to be input to deep neural networks.
Images are generated by padding the full-size images with black pixels from $h\times w$ to $h_{\text{padded}}\times w$, according to
$h_{\text{padded}}=h+\zeta-[(h-w)\bmod\zeta]$ (3)
where $\zeta=100$ is the stride length in pixels. Then, images of dimensions $w\times w$ are cropped every $\zeta$ pixels down the length of the padded image ($h_{\text{padded}}$). The image width $w$ output by the Discover software is $1728$ pixels, so the resulting square cropped images are of dimension $1728\times 1728$. A visualization of this process is shown in Fig. 7. The only other data processing step applied to the images is a normalization around 0 based on the mean and standard deviation of the pixel values of the entire dataset.
Figure 7: Visualization of the strided cropping of the original side scan sonar image to process before inputting into neural networks. Each colored square represents a $w\times w$ crop of the image, with a stride length of $\zeta=100$ px.
@@ 4 Dataset Organization
The dataset is organized by the SSS image and their corresponding labels. We provide the labels as segmentation masks of objects belonging to a shipwreck in the SSS image, including the shipwreck itself and any debris associated with the shipwreck. This includes parts of the shipwreck that may have come off from the shipwreck and parts of the ship that have been dislodged. The images and labels are separated into test and train sets using a 50/50 split. We select a 50/50 split to ensure that the test set is large enough to capture the variance of the shipwrecks sites, ensuring that the robustness of deep learning models can be evaluated through our dataset. In addition, the images are grouped on a per-site basis. This is done to ensure that SSS images that see the same terrain, shipwreck, or debris are not shared between the test and train sets of our dataset.
Table 4: Details of site characteristics, image quality, and number of images per site. Depth refers to the distance from the lake surface to the seafloor at the shipwreck site. Condition is in increasing order with 8 corresponding to most intact. SSS quality is in increasing order with 3 corresponding to highest quality. Terrain corresponds to categorical terrain types as defined in Section 4.2.3. ∗,†,‡ denote shipwrecks that were imaged in a single survey and thus considered as a single “site." Number of images refers to the number of survey legs collected for each site, where each survey leg is processed into one sonar image. Site properties including depth, length, and beam are obtained from Thunder Bay National Marine Sanctuary (2024). We do not present length and beam information for Heart Failure because the wreck is very dispersed. Although the main feature of Artificial Reef is not a wreck, there is one wreck imaged in a single scan at that site, whose length and beam information is unknown. Length and beam information for the Alpena Steamer is also not known.
{tabu}
l|[2pt]c c c c c c c c Name Depth (m) Length (m) Beam (m) Material Condition SSS Quality Terrain # Imgs.
Training
Alpena Steamer $3.4$ – – Wood $4$ $2$ C 10
Bay City $3.4$ $44.5$ $8.8$ Wood $4$ $2$ C 10
D.M. Wilson $12.2$ $54.6$ $9.8$ Wood $6$ $3$ B 22
D.R. Hanna $41.1$ $162.2$ $17.1$ Steel $8$ $2$ A 5
E.B. Allen $30.5$ $40.8$ $7.9$ Wood $6$ $3$ D 24
Egyptian $76.8$ $70.7$ $11.0$ Wood $5$ $2$ D 15
Grecian $30.5$ $90.2$ $12.2$ Steel $7$ $2$ B 5
Harvey Bissell $4.6$ $49.4$ $10.1$ Wood $4$ $2$ C 10
Heart Failure $5.5$ – – Wood $2$ $1$ B 12
Isaac M. Scott $53.3$ $159.7$ $16.5$ Steel $8$ $2$ C 6
Montana $19.2$ $71.9$ $11.0$ Wood $6$ $3$ A 8
Oscar T. Flint $9.1$ $66.4$ $11.3$ Wood $5$ $2$ A 22
Pewabic $50.3$ $61.0$ $9.4$ Wood $6$ $2$ A 5
W.P. Rend $5.2$ $87.5$ $12.2$ Wood $5$ $3$ A 11
Testing
Artificial Reef $3.0$ – – – $1$ $1$ B 6
Barge No. 1 $12.8$ $94.2$ $13.4$ Wood $5$ $2$ A 15
B. Franklin $4.6$ $41.1$ $5.8$ Wood $2$ $1$ B 10
Corsair $55.5$ $40.8$ $7.3$ Wood $5$ $2$ C 4
Corsican $48.8$ $34.1$ $7.6$ Wood $5$ $1$ C 6
James Davidson $10.7$ $30.5$ $9.1$ Wood $4$ $1$ A 10
John F. Warner $2.7$ $38.4$ $7.9$ Wood $3$ $1$ B 6
Haltiner Barge $5.2$ $24.4$ $10.1$ Wood $3$ $2$ A 13
L. van Valkenburg $18.3$ $39.0$ $7.9$ Wood $5$ $3$ B 19
Monohansett $5.5$ $48.8$ $9.1$ Wood $5$ $1$ B 10
Monrovia $42.7$ $136.6$ $17.1$ Steel $8$ $3$ A 8
Shamrock $3.4$ $44.5$ $9.1$ Wood $2$ $1$ B 6
W.H. Gilbert $77.7$ $100.0$ $12.8$ Steel $8$ $3$ C 6
W.P. Thew $25.6$ $40.2$ $7.3$ Wood $5$ $3$ A 17
Viator $57.3$ $70.7$ $10.1$ Steel $8$ $2$ C 11
@@ 4.1 Released Dataset File Structure
In Fig. 8, the released data directory structure is shown. The data is split into test and train, according to our iterative approach as described in this paper. Within the test and train directories, the sonar images and their corresponding binary labels are stored. Each sonar image (.PNG format) corresponds to one leg of a survey mission. The site of the survey is indicated in the image file names. We also include a directory with four sites of just surveyed terrain in the extras directory. The raw .JSF sonar files – which contain AUV GPS and positioning data – will be released on the NOAA National Centers for Environmental Information database and linked to the project webpage.
4.2 Train-Test Split Selection
Table 4 provides details of each site, organized by train and test split. Our goal with the train-test split is to ensure that each set provides enough diversity of data for both the training and evaluation of deep learning models. We inform our split selection through both expert-informed and data-driven metrics, with the target of capturing the underlying distribution of the data as uniformly as possible. The atomic unit we use to split up the dataset into train and test sets is each survey site location. This was to ensure that no sonar images overlapping the same region are present in both the test and train sets. Please note in Table 4 some sites contain multiple shipwrecks, as indicated by the superscripts. To inform the train-test split selection, we group the sites based on three major categories: shipwreck condition, sonar image quality, and terrain type. For each site, a vector score of length three is assigned based on these quantitative and qualitative metrics. A 50/50 train-test split is then performed using an iterative stratification on our multi-labeled data (Sechidis et al., 2011).
@@ 4.2.1 Wreck Condition:
We leverage domain expert knowledge from scientists at TBNMS in order to assign wreck condition labels. Experts who have conducted dive surveys to inspect these ships provided a scale of 1-8 for wreck condition. There are five sites labeled 8 for complete and intact, one site labeled 7 for complete but collapsed, four sites labeled 6 for semi-complete, nine sites label 5 for partially intact, four sites labeled 4 for fragmented, two sites labeled 3 for fragmented and disarticulated, three sites labeled 2 for widely dispersed debris fields, and one site labeled 1 for non-shipwreck cultural materials.
Table 6: Per-site IoUship compared across each baseline for each site. Higher is better ($\uparrow$).
{tabu}
c|[2pt]cccccccccccccc Artificial Reef Barge No. 1 Corsair Corsican Davidson W.H. Gilbert Haltiner Barge L. Van Valkenburg Monohansett Monrovia Shamrock W.P. Thew Viator
Yang et. al 0.0020.4710.2310.0560.0000.2760.0310.3670.0200.4780.1020.1610.563
ViT-Adapter0.0030.4120.5420.0880.0000.0060.3680.1270.4960.4690.3540.1760.631
DeepLabv30.0030.6700.5310.0420.0310.6580.4590.6450.0340.4110.1510.4190.667
HRNet0.0110.7030.2930.1740.0670.6410.4480.6410.2070.5660.0030.4370.646
UNet0.0060.7360.5640.1160.0740.7260.2420.6410.0420.5720.4280.5450.646
SOD0.0170.7750.5830.0770.0390.7490.4420.7130.4670.5720.0010.5800.776
@@ 4.2.2 Sonar Image Quality Assessment:
The sonar image quality assessment is done based on well-developed metrics for optical image quality assessment (IQA). We utilize the no-reference, completely blind IQA metric Natural Image Quality Evaluator (NIQE), which constructs a quality score from a natural scene statistic model trained on undistorted images (Mittal et al., 2013). Although there are some sonar quality metrics (Liu et al., 2023; Chen et al., 2020), they generally rely on reference images and measure signal degradation during transmission, which is not suitable for our dataset. There are many challenges in evaluating the quality of a sonar image, the first of which being that many IQA metrics are targeted at measuring quality deterioration from compression. However, our sonar images are not compressed – rather, any distortion in the images is a result of when the Iver3 is surveying shallower sites, making it more vulnerable to drift from waves (Fig. 6d). If the shipwreck partially falls into the nadir, this can cause significant interference and acoustic shadows that disrupt the geometry of the wreck (Fig. 6e). Additionally, if the shipwreck falls into the outer edges of the sonar image, the increased noise of more distant acoustic returns causes significant noise in the final sonar image (Fig. 6f). Informed by the NIQE score, we assign classes of 1-3 to the images where 1 indicates poorer sonar image quality and 3 indicates higher image quality.